Students in a math class need an average of at least 90 points to earn an A. One student's test scores are 88, 91, and 85. What must the student score on the next test to earn an A? Show all work and steps. Use an equation.
Answer by checkley77(12844)

(88+91+85+x)/4>=90
(264+x)/4>=90
264+x>=4*90
264x>=360
x>=360-264
x>=96 is the minimum score on the final test.